# Step-Aware Residual-Guided Diffusion for EEG Spatial Super-Resolution

### v1

@@ Abstract
For real-world BCI applications, lightweight Electroencephalography (EEG) systems offer the best cost–deployment balance. However, such spatial sparsity of EEG limits spatial fidelity, hurting learning and introducing bias. EEG spatial super-resolution methods aim to recover high-density EEG signals from sparse measurements, yet is often hindered by distribution shift and signal distortion and thus reducing fidelity and usability for EEG analysis and visualization. To overcome these challenges, we introduce SRGDiff, a step-aware residual-guided diffusion model that formulates EEG spatial super-resolution as dynamic conditional generation. Our key idea is to learn a dynamic residual condition from the low-density input that predicts the step-wise temporal and spatial details to add and uses the evolving cue to steer the denoising process toward high density reconstructions. At each denoising step, the proposed residual condition is additively fused with the previous denoiser feature maps, then a step-dependent affine modulation scales and shifts the activation to produce the current features. This iterative procedure dynamically extracts step-wise temporal rhythms and spatial-topographic cues to steer high-density recovery and maintain a fidelity–consistency balance. We adopt a comprehensive evaluation protocol spanning signal-, feature-, and downstream-level metrics across SEED, SEED-IV, and Localize-MI and multiple upsampling scales. SRGDiff achieves consistent gains of up to 40% over strong baselines, proving its superiority in the task of EEG spatial super-resolution. Moreover, topographic visualizations comparison and substantial EEG-FID gains jointly indicate that our SR EEG mitigates the spatial–spectral shift between low- and high-density recordings. Our code is available at https://anonymous.4open.science/r/DDPM-VAE-6F6B.
@@ 1 Introduction
Electroencephalography (EEG) is a noninvasive technique for monitoring the brain’s electrical activity, with widespread applications in neuroscience and clinical practice—ranging from brain–computer interfaces and epilepsy diagnosis to emotion recognition Jiang et al. (2025). However, EEG’s spatial resolution is inherently constrained by the number of scalp electrodes and the volume‐conduction effect Li et al. (2025a). High‐density (HD) systems with hundreds of channels can mitigate these issues but are costly, cumbersome to deploy, and uncomfortable for extended wear, whereas low‐density (LD) setups (e.g., 8 or 16 electrodes) are far more practical yet suffer from severe under‐sampling bias Wang et al. (2025). Indeed, as illustrated in Figure 1(c), the inter‐channel activation patterns of 256‐channel HD EEG diverge dramatically from those of 16‐channel LD EEG, highlighting the strong bias in sparse recordings. EEG spatial super-resolution (SR) has therefore garnered growing attention, with methods that reconstruct high-density EEG from sparse recordings increasingly explored and applied.
Traditionally, EEG spatial super‐resolution has relied on direct feature‐mapping techniques that learn an end-to-end mapping from low-density to high-density representations. These methods fall into two main categories: one employs convolutional neural networks or Transformers to upsample LD feature maps into HD ones Tang et al. (2022), and the other leverages generative adversarial networks-based architectures that synthesize SR EEG signals conditioned on LD inputs Wang et al. (2024). However, by treating the mapping as a static projection, these approaches often oversimplify the complex, nonlinear inter-electrode dependencies or demand vast amounts of training data and compute, resulting in overly smooth, detail-poor reconstructions that fail to capture true spatial consistency. Figure 1(c) indicates that such feature-mapping methods merely extend LD information, rather than recovering authentic HD channel relationships.
@@
Recently, diffusion models have been widely applied to time-series generation and missing-data imputation Huang et al. (2025); Yuan & Qiao (2024); Li et al. (2025b). In this context, EEG spatial super-resolution can be cast as conditional generation, where LD observations guide the recovery of HD signals. Within this line of work, researchers mainly focused on conditioning strategies through concatenating low-density features with the noise input Vetter et al. (2024) or using cross-attention between modalities Wang et al. (2025) as shown in Figure 1(a). While effective in practice, these approaches remain susceptible to a consistency–fidelity trade-off. Interpolation-oriented SR tends to cause distribution shift, making reconstructions adhere too closely to the LD observation and deviate from the HD ground truth. Conversely, generation-oriented SR often introduces distortion, producing HD-like content that fails to remain consistent with the LD input.
@@
  • We propose a dynamic residual guidance paradigm: the path residual estimated from LD inputs serves as a per-step directional correction and is fused additively for incremental updates, yielding a stable, step-aware sampling scheme transferable to general SR/MDI settings.
@@ 2 Related Work
Diffusion‐based models have emerged as a powerful framework for time‐series imputation, leveraging denoising diffusion processes to reconstruct missing values. Diffusion-TS Yuan & Qiao (2024) demonstrates interpretable conditional generation across diverse time‐series without domain‐specific priors. More recently, RDPI Liu et al. (2025) further enhances precision and efficiency by first generating coarse estimates of missing values through deterministic interpolation, then conditioning a diffusion model on both observed data and these estimates to iteratively refine residual errors. SaSDim Zhang et al. (2024) introduces self‑adaptive noise scaling to preserve spatial dependencies within sensor networks, while SADI Islam et al. (2025) integrate self‑attention mechanisms to handle partial data missing.
@@
Early attempts at EEG spatial super‐resolution adapted image‐based frameworks to reconstruct dense electrode maps from sparse recordings. EEGSR‐GAN Corley & Huang (2018) first applied adversarial training to hallucinate missing channels. ESTformer Li et al. (2025a) then introduced spatiotemporal transformers to model long‐range dependencies across electrodes and successfully capture global patterns. More recent diffusion‐based and attention‐driven approaches have sought to address these limitations. DDPM‐EEGVetter et al. (2024) leverages denoising diffusion probabilistic models to iteratively refine spatial patterns. STAD Wang et al. (2025) tackles this by decomposing spatial–temporal interactions into spatial-temporal attention streams. This diffusion-based generative paradigm improves diversity and spectral fidelity compared to GANs, yet their static condition can still lead to distribution drifts and distortion.
@@ 4.1 Latent Diffusion Model Backbone
$\mathcal{L}_{\text{VAE}}=\|\widehat{X}^{H}-X^{H}\|_{2}^{2}+\lambda_{\text{spec}}\big\|\mathrm{STFT}(\widehat{X}^{H})-\mathrm{STFT}(X^{H})\big\|_{1}+\lambda_{KL}\,\mathrm{KL}\!\left(q_{E}(z\!\mid\!X^{H})\,\|\,\mathcal{N}(0,I)\right).$ (5)
@@ 4.3 Step-Aware Modulation Module
$\widetilde{h}_{t}=\sigma_{t}h_{t}+(1-\sigma_{t})e_{t}=\gamma_{t}E_{SMM}(c)+(1-\gamma_{t})e_{t}.$ (9)
@@ 4.4 Training Strategy
$\mathcal{L}_{\text{Stage 2}}=\mathbb{E}_{z_{0},\epsilon,t}\big[\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\|_{2}^{2}\big]+\lambda_{res}\sum_{t=1}^{T}\|R_{\varphi}(c,t)-(z_{t}-z_{0})\|_{2}^{2}+\lambda_{SMM}(\|\gamma_{t}-1\|_{2}^{2}+\|\beta_{t}\|_{2}^{2}).$ (11)
@@ 5.1 Downstream Datasets
In this study, we employ three publicly available EEG datasets. The SEED dataset Zheng & Lu (2015) uses $15$ film clips of approximately four minutes each as emotional stimuli to induce stable and continuous emotional responses in three categories: positive, neutral, and negative. Data were acquired via $62$ channels at $1000$ Hz, downsampled to $200$ Hz, band-pass filtered (07575 Hz) and segments with faulty sensors removed. SEED-IV Zheng et al. (2018) extends SEED by using the same $15$ subjects and $62$-channel setup ($1000$ Hz) but adds music and image stimuli to evoke happiness, sadness, fear and neutral states; preprocessing mirrors that of SEED. Localize-MI Mikulan et al. (2020) contains 61 presurgical sessions from seven drug-resistant epilepsy patients, where $256$-channel scalp EEG was recorded at $8000$ Hz during 0.10.155 mA intracerebral single-pulse stimulation; preprocessing includes a $0.1$ Hz high-pass filter, notch filter, bad-channel/trial removal and alignment of trials to the -$300$ ms to +$50$ ms stimulus-artifact window.
@@ 5.2 Experiment Setup
We compare SRGDiff with strong EEG SR and time-series imputation baselines: ESTformer Li et al. (2025a) and STAD Wang et al. (2025) (transformer-/diffusion-based EEG SR), DDPMEEG Vetter et al. (2024) (diffusion for ECoG SR), SaSDim Zhang et al. (2024) and SADI Islam et al. (2025) (advanced missing data imputation), and the two-stage residual method RDPI Liu et al. (2025). We use authors’ official implementations when available and otherwise provide carefully verified reimplementations, applying their recommended hyperparameters and unifying training epochs and sampling steps across methods.
@@
Evaluation Protocol.
@@ 5.3 Main Results
We also report EEG-FID results in Figure 3, and our method consistently achieves the lowest FID scores across SEED, SEED-IV, and Localize-MI datasets under different scale factors. These results indicate that our approach generates EEG signals that are statistically closer to the real distribution in the temporal domain.
@@
5.4 Topology Visualization Results
We further analyze the spectral fidelity of generated signals by visualizing EEG topographic maps under different scale factors. An EEG topographic map projects the power spectral density (PSD) of each channel onto the scalp surface, providing an intuitive representation of the spatial distribution of oscillatory energy. As shown in Figure 4, although our reconstructed signals still exhibit minor differences from the original data, they preserve a high degree of overlap in critical regions with strong PSD responses.
@@ 5.5 Downstream tasks
Method SEED SEED-IV MI
@@ 5.6 Ablation Study
To evaluate the contribution of each module in SRGDiff to EEG super-resolution reconstruction, we conducted ablation studies comparing SRGDiff with three variant models. LDM retains only the pre-trained VAE and DDIM; LDM+SMM preserves the Step-aware modulation module; LDM+RDM retains the Residual Direction Module. All models were tested under the same experimental settings.
@@
Figure 5 reports NMSE and PCC across $2\times$, $4\times$ and $8\times$ upsampling in SEED dataset. At $8\times$, adding SMM to the baseline LDM cuts NMSE from $0.86$ to $0.45$ with 4747% reduction and boosts PCC from $0.34$ to $0.69$, demonstrating its effectiveness in temporally aligning the denoising trajectory. Incorporating RDM yields a comparable NMSE reduction with 4444% and raises PCC to $0.67$, highlighting its role in injecting prior information for spatial consistency. When combined in SRGDiff, these modules further decrease NMSE to $0.34$ with 6060% overall reduction and elevate PCC to $0.81$. More ablation results in SEED-IV and Localize-MI datasets are shown in Figure 10 in the Appendix.
@@ B.1 Dataset Details
Localize-MI (available at https://doi.org/10.12751/g-node.1cc1ae) is a high-density intracranial EEG dataset from seven drug-resistant epilepsy patients during $61$ presurgical sessions. Stereo–EEG electrodes delivered single-pulse biphasic currents (0.150.1–5 mA), and $256$ channels were recorded at $8000$ Hz. Preprocessing included $0.1$ Hz high-pass filtering, notch filters at 5050/100100/150150/$200$ Hz, bad-channel/trial removal, and trial alignment using stimulation artifact peaks (300–300 to +50+50 ms window). In the Localize-MI dataset, we designed a binary classification task (epileptic vs. nonepileptic) to evaluate the effectiveness of synthetic super-resolution EEG (SR EEG) in detecting epileptic abnormalities. Specifically, EEG signals recorded before electrical stimulation are labeled as nonepileptic, while those recorded during stimulation are labeled as epileptic. The detailed experimental setup follows the description provided in the STAD Wang et al. (2025) model section.

### v2

@@ Abstract
For real-world brain–computer interface (BCI) applications, lightweight Electroencephalography (EEG) systems offer the best cost–deployment balance. However, such spatial sparsity of EEG limits spatial fidelity, hurting learning and introducing bias. EEG spatial super-resolution methods aim to recover high-density EEG signals from sparse measurements, yet is often hindered by distribution shift and signal distortion and thus reducing fidelity and usability for EEG analysis and visualization. To overcome these challenges, we introduce SRGDiff, a step-aware residual-guided diffusion model that formulates EEG spatial super-resolution as dynamic conditional generation. Our key idea is to learn a dynamic residual condition from the low-density input that predicts the step-wise temporal and spatial details to add and uses the evolving cue to steer the denoising process toward high density reconstructions. At each denoising step, the proposed residual condition is additively fused with the previous denoiser feature maps, then a step-dependent affine modulation scales and shifts the activation to produce the current features. This iterative procedure dynamically extracts step-wise temporal rhythms and spatial-topographic cues to steer high-density recovery and maintain a fidelity–consistency balance. We adopt a comprehensive evaluation protocol spanning signal-, feature-, and downstream-level metrics across SEED, SEED-IV, and Localize-MI and multiple upsampling scales. SRGDiff consistently achieves higher SNR than the baseline ESTformer and STAD among Localize-MI, SEED and SEED-IV datasets, with up to roughly $75\%$ relative SNR improvement in the most challenging $8\times$ setting. Moreover, topographic visualizations comparison and substantial EEG-FID gains jointly indicate that our SR EEG mitigates the spatial–spectral shift between low- and high-density recordings. Our code is available at https://github.com/DhrLhj/ICLR2026SRGDiff.
@@ 1 Introduction
Electroencephalography (EEG) is a noninvasive technique for monitoring the brain’s electrical activity, with widespread applications in neuroscience and clinical practice—ranging from brain–computer interfaces and epilepsy diagnosis to emotion recognition (Jiang et al., 2025). However, EEG’s spatial resolution is inherently constrained by the number of scalp electrodes and the volume‐conduction effect (Li et al., 2025a). High‐density (HD) systems with hundreds of channels can mitigate these issues but are costly, cumbersome to deploy, and uncomfortable for extended wear, whereas low‐density (LD) setups (e.g., 8 or 16 electrodes) are far more practical yet suffer from severe under‐sampling bias (Wang et al., 2025). Indeed, as illustrated in Figure 1(c), the inter‐channel activation patterns of 256‐channel HD EEG diverge dramatically from those of 16‐channel LD EEG, highlighting the strong bias in sparse recordings. EEG spatial super-resolution (SR) has therefore garnered growing attention, with methods that reconstruct high-density EEG from sparse recordings increasingly explored and applied.
Traditionally, EEG spatial super‐resolution has relied on direct feature‐mapping techniques that learn an end-to-end mapping from low-density to high-density representations. These methods fall into two main categories: one employs convolutional neural networks or Transformers to upsample LD feature maps into HD ones (Tang et al., 2022), and the other leverages generative adversarial networks-based architectures that synthesize SR EEG signals conditioned on LD inputs (Wang et al., 2024). However, by treating the mapping as a static projection, these approaches often oversimplify the complex, nonlinear inter-electrode dependencies or demand vast amounts of training data and compute, resulting in overly smooth, detail-poor reconstructions that fail to capture true spatial consistency. Figure 1(c) indicates that such feature-mapping methods merely extend LD information, rather than recovering authentic HD channel relationships.
@@
Recently, diffusion models have been widely applied to time-series generation and missing-data imputation (Huang et al., 2025; Yuan and Qiao, 2024; Li et al., 2025b). In this context, EEG spatial super-resolution can be cast as conditional generation, where LD observations guide the recovery of HD signals. Within this line of work, researchers mainly focused on conditioning strategies through concatenating low-density features with the noise input (Vetter et al., 2024) or using cross-attention between modalities (Wang et al., 2025) as shown in Figure 1(a). While effective in practice, these approaches remain susceptible to a consistency–fidelity trade-off. Interpolation-oriented SR tends to cause distribution shift, making reconstructions adhere too closely to the LD observation and deviate from the HD ground truth. Conversely, generation-oriented SR often introduces distortion, producing HD-like content that fails to remain consistent with the LD input.
@@
  • We propose a dynamic residual guidance paradigm: the path residual estimated from LD inputs serves as a per-step directional correction and is fused additively for incremental updates, yielding a stable, step-aware sampling scheme that remains effective across datasets and a wide range of SR factors.
@@ 2 Related Work
Diffusion‐based models have emerged as a powerful framework for time‐series imputation, leveraging denoising diffusion processes to reconstruct missing values. Diffusion-TS (Yuan and Qiao, 2024) demonstrates interpretable conditional generation across diverse time‐series without domain‐specific priors. More recently, RDPI (Liu et al., 2025) further enhances precision and efficiency by first generating coarse estimates of missing values through deterministic interpolation, then conditioning a diffusion model on both observed data and these estimates to iteratively refine residual errors. SaSDim (Zhang et al., 2024) introduces self‑adaptive noise scaling to preserve spatial dependencies within sensor networks, while SADI (Islam et al., 2025) integrate self‑attention mechanisms to handle partial data missing.
@@
Early attempts at EEG spatial super‐resolution adapted image‐based frameworks to reconstruct dense electrode maps from sparse recordings. EEGSR‐GAN (Corley and Huang, 2018) first applied adversarial training to hallucinate missing channels. ESTformer (Li et al., 2025a) then introduced spatiotemporal transformers to model long‐range dependencies across electrodes and successfully capture global patterns. More recent diffusion‐based and attention‐driven approaches have sought to address these limitations. DDPM‐EEG (Vetter et al., 2024) leverages denoising diffusion probabilistic models to iteratively refine spatial patterns. STAD (Wang et al., 2025) tackles this by decomposing spatial–temporal interactions into spatial-temporal attention streams. This diffusion-based generative paradigm improves diversity and spectral fidelity compared to GANs, yet their static condition can still lead to distribution drifts and distortion.
@@ 4.1 Latent Diffusion Model Backbone
$\mathcal{L}_{\text{VAE}}=\|\widehat{X}^{H}-X^{H}\|_{2}^{2}+\lambda_{\text{spec}}\big\|\mathrm{STFT}(\widehat{X}^{H})-\mathrm{STFT}(X^{H})\big\|_{1}+\lambda_{KL}\,\mathrm{KL}\!\left(q_{E}(z\!\mid\!X^{H})\,\|\,\mathcal{N}(0,I)\right),$ (5)
where $\mathrm{STFT}(\cdot)$ denotes the short-time Fourier transform applied along the temporal dimension of each EEG channel to encourage spectral fidelity.
@@ 4.3 Step-Aware Modulation Module
$\widetilde{h}_{t}=\sigma_{t}h_{t}+(1-\sigma_{t})e_{t}=\sigma_{t}E_{SMM}(c)+(1-\sigma_{t})e_{t}.$ (9)
@@ 4.4 Training Strategy
$\mathcal{L}_{\text{Stage 2}}=\mathbb{E}_{z_{0},\epsilon,t}\big[\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\|_{2}^{2}\big]+\lambda_{res}\sum_{t=1}^{T}\|R_{\varphi}(c,t)-(z_{0}-z_{t})\|_{2}^{2}+\lambda_{SMM}(\|\gamma_{t}-1\|_{2}^{2}+\|\beta_{t}\|_{2}^{2}).$ (11)
@@ 5.1 Downstream Datasets
In this study, we employ three publicly available EEG datasets. The SEED dataset (Zheng and Lu, 2015) uses $15$ film clips of approximately four minutes each as emotional stimuli to induce stable and continuous emotional responses in three categories: positive, neutral, and negative. Data were acquired via $62$ channels at $1000$ Hz, downsampled to $200$ Hz, band-pass filtered (07575 Hz) and segments with faulty sensors removed. SEED-IV (Zheng et al., 2018) extends SEED by using the same $15$ subjects and $62$-channel setup ($1000$ Hz) but adds music and image stimuli to evoke happiness, sadness, fear and neutral states; preprocessing mirrors that of SEED. Localize-MI (Mikulan et al., 2020) contains 61 presurgical sessions from seven drug-resistant epilepsy patients, where $256$-channel scalp EEG was recorded at $8000$ Hz during 0.10.155 mA intracerebral single-pulse stimulation; preprocessing includes a $0.1$ Hz high-pass filter, notch filter, bad-channel/trial removal and alignment of trials to the -$300$ ms to +$50$ ms stimulus-artifact window.
@@ 5.2 Experiment Setup
We compare SRGDiff with strong EEG SR and time-series imputation baselines: ESTformer (Li et al., 2025a) and STAD (Wang et al., 2025) (transformer-/diffusion-based EEG SR), DDPMEEG (Vetter et al., 2024) (diffusion for ECoG SR), SaSDim (Zhang et al., 2024) and SADI (Islam et al., 2025) (advanced missing data imputation), and the two-stage residual method RDPI (Liu et al., 2025). We use authors’ official implementations when available and otherwise provide carefully verified reimplementations, applying their recommended hyperparameters and unifying training epochs and sampling steps across methods.
@@
5.3 Evaluation Protocol.
We assess SR quality at three complementary levels to balance faithfulness to the ground truth, preservation of neurophysiological structure, and practical utility.
Signal level
(does the waveform match?): We follow ESTformer and report normalized mean squared error (NMSE), Pearson correlation coefficient (PCC), and reconstruction signal-to-noise ratio (SNR) with respect to the HD reference, plus topology maps for qualitative inspection (formal definitions in the Appendix).
Feature level
(does the representation distribution match?): We adopt EEG-FID following Lai et al. (2025), using a frozen EEGNet trained per dataset on its training split; the embedding dimension is 256 for SEED/SEED-IV and 512 for Localize-MI. In addition, we report a frequency-domain MAE: we first compute channel-wise STFTs of the reconstructed and reference HD EEG, form their power spectra, and then compute the normalized mean squared error between these spectra averaged over channels and frequency bins, so as to capture spectral distortions that are not reflected by time-domain NMSE alone.
Downstream level
@@ 5.4 Main Results
5.5 Feature-level Reconstruction Evaluation
We report EEG-FID results in Figure 3, and our method consistently achieves the lowest FID scores across SEED, SEED-IV, and Localize-MI datasets under different scale factors. These results indicate that our approach generates EEG signals that are statistically closer to the real distribution in the temporal domain.
We further analyze the spectral fidelity of generated signals by visualizing EEG topographic maps under different scale factors. An EEG topographic map projects the power spectral density (PSD) of each channel onto the scalp surface, providing an intuitive representation of the spatial distribution of oscillatory energy. As shown in Figure 4 and Figure  15, although our reconstructed signals still exhibit minor deviations from the original data, they preserve a high degree of overlap in critical regions with strong PSD responses.
@@ 5.5 Feature-level Reconstruction Evaluation
Beyond qualitative inspection, we also report a frequency-domain error metric that quantifies the mean absolute error between reconstructed and real HD topomaps. As shown in Table 3, SRGDiff consistently achieves the lowest frequency-domain MAE across datasets and SR scales, indicating better preservation of the spatial distribution of spectral power.
Model SEED SEED-IV Localize-MI
[…]
ESTformer 6.96 9.31 9.73 7.11 8.30 8.86 7.03 16.67 32.32 35.73
STAD 9.19 11.04 14.40 9.50 10.95 13.12 8.76 13.11 22.53 25.37
SRGDiff 3.89 5.12 4.95 3.99 4.08 4.84 3.86 7.30 11.50 13.76
Table 3: Frequency-domain MAE between reconstructed and real HD topomaps on SEED, SEED-IV, and Localize-MI under different SR factors.
@@ 5.6 Downstream tasks
Method SEED SEED-IV Localize-MI
@@ 5.7 Ablation Study
To evaluate the contribution of each module in SRGDiff to EEG super-resolution reconstruction, we conducted ablation studies comparing SRGDiff with three variant models. LDM+LD keeps only the VAE–DDIM backbone and takes the LD EEG as its input condition; LDM+SMM preserves the Step-aware modulation module; LDM+RDM retains the Residual Direction Module. All models were tested under the same experimental settings.
@@
Figure 5 reports NMSE, PCC and SNR across $2\times$, $4\times$ and $8\times$ upsampling in SEED dataset. At $8\times$, adding SMM to the baseline cuts NMSE from $0.86$ to $0.45$ with 4747% reduction and boosts PCC from $0.34$ to $0.69$, demonstrating its effectiveness in temporally aligning the denoising trajectory. Incorporating RDM yields a comparable NMSE reduction with 4444% and raises PCC to $0.67$, highlighting its role in injecting prior information for spatial consistency. When combined in SRGDiff, these modules further decrease NMSE to $0.34$ with 6060% overall reduction and elevate PCC to $0.81$. More ablation results in SEED-IV and Localize-MI datasets are shown in Figure 11 in the Appendix.
@@ 6 Conclusion
7 Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2022ZD0118001), the National Natural Science Foundation of China (Grant Nos. U22A2022, 62332017), and was conducted in part at the MOE Key Laboratory of Advanced Materials and Devices for Post-Moore Chips, the Beijing Key Laboratory of Big Data Innovation and Application for Skeletal Health Medical Care, and the Beijing Advanced Innovation Center for Materials Genome Engineering.
@@ B.1 Dataset Details
Localize-MI (available at https://doi.org/10.12751/g-node.1cc1ae) is a high-density intracranial EEG dataset from seven drug-resistant epilepsy patients during $61$ presurgical sessions. Stereo–EEG electrodes delivered single-pulse biphasic currents (0.150.1–5 mA), and $256$ channels were recorded at $8000$ Hz. Preprocessing included $0.1$ Hz high-pass filtering, notch filters at 5050/100100/150150/$200$ Hz, bad-channel/trial removal, and trial alignment using stimulation artifact peaks (300–300 to +50+50 ms window). In the Localize-MI dataset, we designed a binary classification task (epileptic vs. nonepileptic) to evaluate the effectiveness of synthetic super-resolution EEG (SR EEG) in detecting epileptic abnormalities. Specifically, EEG signals recorded before electrical stimulation are labeled as nonepileptic, while those recorded during stimulation are labeled as epileptic. The detailed experimental setup follows the description provided in the STAD (Wang et al., 2025) model section.